OpenStudy (helpconfusedami):

PPLLEEAASE HELP! Very urgent for a test tomorrow. May someone help me understand circuits basic physics? Such as interpreting circuit diagrams for resistance, voltage, and current! I will give you a thumbs up! :)

OpenStudy (bmk614):

Know Ohm's Law is V=IR To find the equivalent resistance of series circuits \[R_{eq}=R_1+R_2+...\] To find the equivalent resistance of parallel circuits \[\frac{ 1 }{R_{eq}? }=\frac{ 1 }{ R_1 }+\frac{ 1 }{ R_2 }+...\]

OpenStudy (bmk614):

You need to add up all the resistors to find the equivalent resistance of the whole thing. Then divide the Voltage by the Resistance to find the Amps. Amps stay the same in series. Voltage stays the same in parallel.

OpenStudy (helpconfusedami):

hmm... what if there is an open switch?

OpenStudy (helpconfusedami):

for example, if there is a battery then 3 lightbulbs in a circuit, and the switch is not allowing electrons to move through the wire-

OpenStudy (bmk614):

The open switch just means that there is no current going through it.

OpenStudy (helpconfusedami):

is the batterie going to have no current ?

OpenStudy (helpconfusedami):

okay..

OpenStudy (helpconfusedami):

will the circuit have voltage? and will the total resistance be the same?

OpenStudy (bmk614):

The battery still has the same voltage.

OpenStudy (helpconfusedami):

Okay. Do the batteries have any voltage?

OpenStudy (helpconfusedami):

*i mean lights Sorry.

OpenStudy (bmk614):

No because they only have one charge.